# Kudla Millson lift of toric cycles and restriction of Hilbert modular forms

### v3

@@ 1.1.  Intersection numbers of geodesics on modular curves
Let $Y_{0}(p)=\Gamma_{0}(p)\backslash\mathscr{H}$ be the open modular curve for some prime $p$. Let $\gamma_{\infty}$ be the image in $Y_{0}(p)$ of the geodesic in $\mathscr{H}$ from 00 to $\infty$. It defines a relative cycle
@@
If two geodesics $\gamma_{1}$ and $\gamma_{2}$ in $Y_{0}(p)$ intersect transversely and in a compact set then $\langle\gamma_{1},\gamma_{2}\rangle=\sum_{z\in\gamma_{1}\cap\gamma_{2}}\pm 1$ is the the topological intersection number, where ±$\pm 1$ depends on the local orientation at the intersection point. Darmon-Pozzi-Vonk prove the following in [DPV21, Theorem. A].
@@
of weight $2$ and level $\Gamma_{0}(p)$ is the diagonal restriction of a $p$-stabilized Hilbert-Eisenstein series $E^{(p)}(\psi,\tau,\tau)$ for $\operatorname{SL}_{2}(\mathbb{Q}(\sqrt{D}))$.
In [Bra23], we showed how to recover (and generalize) the theorem of Darmon-Pozzi-Vonk by using the Kudla-Millson lift. The idea was to consider the embedding $F^{\times}\subset\operatorname{SO}(d,d)$ of a totally real field $F$ of degree $d$, which gives a non-compact cycle on the locally symmetric space of $\operatorname{SO}(d,d)$. The (regularized) integral of the Kudla-Millson theta lift over this cycles gives a generating series of intersection numbers generalizing the right handside of (1.4). On the other hand, the relation to a Hilbert-Eisenstein series follows from a Siegel-Weil formula and the seesaw
[…]
In the case where $F$ is real quadratic field, this yields to the theorem of Darmon-Pozzi-Vonk above.
@@
Another closely related seesaw appears in [BCG20]
[…]
where the right hand side is the diagonal restriction of (the same) Hilbert-Eisenstein series, and the left hand side is the evaluation of an Eisenstein class on the same cycle above. The two seesaws are related by embedding $\operatorname{GL}_{d}(\mathbb{Q})$ in $\operatorname{SO}(d,d)$, and the Bergeron-Charollois-Garcia lift is closely related to the Kudla-Millson lift via the Mathai-Quillen form. We hope to explain this in more detail in future work.
@@ 1.2.  Main result
In this paper, we replace the torus $F^{\times}\subset\operatorname{SO}(d,d)$ by a more general anisotropic maximal $\mathbb{Q}$-torus, of maximal $\mathbb{R}$-rank in an orthogonal group $\operatorname{SO}(p,q)$ of signature $(p,q)$ with $p\geq q>0$. We restrict ourselves to anisotropic tori, to avoid having to deal with the regularization of the integral as in [Bra23]. However, one can combine the result presented here with the result of loc. cit. to consider any maximal $\mathbb{Q}$-torus of maximal $\mathbb{R}$-rank, not necessarily anisotropic.
@@
The study of cycles attached to algebraic tori in orthogonal groups is also motivated by the recent work of Darmon-Gehrmann-Linowski. In [DGL23], the authors define rigid meromorphic cocycles in arbitrary signature, generalizing the rigid meromorphic cocycles in signature $(2,2)$ studied in [DPV21]. They conjecture that evaluated on toric cycles, these cocycles should be algebraic. The seesaws presented here could be relevant.
@@
Let $L\subset V$ be an even integral lattice of level $N$. Let $K_{f}\subset\operatorname{SO}_{V}(\widehat{\mathbb{Q}})$ be an open compact subgroup stabilizing the finite Schwartz function $\varphi=\mathbf{1}_{\widehat{L}}\in\mathcal{S}(V_{\widehat{\mathbb{Q}}})$, where $\widehat{L}=L\otimes\widehat{\mathbb{Z}}\subset V_{\widehat{\mathbb{Q}}}$ and $\widehat{\mathbb{Q}}$ are the finite adèles. We set $K=K_{\infty}K_{f}$ and consider the adelic space
@@
In [KM86, KM87], Kudla and Millson construct an element
$\Theta_{\varphi}\in H^{q}(Y)\otimes M_{d}(\Gamma_{0}(N))$ (1.6)
that realizes a lift from the homology to the space of modular forms of weight $d=\frac{p+q}{2}$ and level $\Gamma_{0}(N)$. As for the modular curve, there is a homological pairing
$\langle-,-\rangle\colon H_{q}(Y,\mathbb{Z})\times H_{pq-q}(Y,\mathbb{Z})\longrightarrow\mathbb{Z}.$ (1.7)
If the homology classes are represented by two smooth immersed submanifolds that intersect transversely and in a compact set, then and the intersection number $\langle C_{1},C_{2}\rangle$ is the signed intersection number, as for geodesics. The main feature of the Kudla-Millson lift is that for a cycle $C$ in $H_{q}(Y)$ it has the Fourier expansion
$\Theta_{\varphi}(\tau,C)\coloneqq\int_{C}\Theta_{\varphi}(\tau)=c^{0}_{\varphi}(C)+\sum_{n=1}^{\infty}\langle C,C_{n}(\varphi)\rangle q^{n},$ (1.8)
@@
$\chi=(\chi_{\infty},\chi_{f})\colon\operatorname{T}(\mathbb{Q})\backslash\operatorname{T}(\mathbb{A})\longrightarrow\mathbb{C}^{\times}$
@@
$\mathcal{C}_{T}\coloneqq\operatorname{T}(\mathbb{Q})^{+}\backslash\operatorname{T}(\widehat{\mathbb{Q}})/K_{T,f}.$ (1.9)
@@
where $\Lambda\coloneqq\operatorname{T}(\mathbb{Q})^{+}\cap K_{T,f}$. The embedding of $\operatorname{T}(\mathbb{A})$ in $\operatorname{SO}_{V}(\mathbb{A})$ induces an immersion $Y_{T}\longrightarrow Y$, and the image of the connected component $\Lambda\backslash\mathbb{R}_{>0}^{q}$ associated to $I\in\mathcal{C}_{T}$ defines a cycle $C_{T,I}\in\mathcal{Z}_{q}(Y)$. Let us now suppose that $\chi_{f}$ is trivial on $K_{T,f}$. We can see the finite part $\chi_{f}$ as a function on $\chi_{f}\colon\mathcal{C}_{T}\longrightarrow\mathbb{C}^{\times}$ and define the cycle
@@
By the work of Kudla-Millson, it is known that if $q$ is odd, then $\Theta_{\varphi}(\tau,C)$ is a cusp form for any cycle $C$; see [KM90, Theorem. 2]. We show that when we restrict to the cycles $C_{\chi}$, then the lift is always a cusp form.
@@
Let $V$ be an even dimensional quadratic space of signature $(p,q)$ with $p\geq q>0$ and let $C_{\chi}$ be the cycle attached to an anisotropic maximal $\mathbb{Q}$-torus of maximal real rank $q$. Then $\Theta_{\varphi}(\tau,C_{\chi})\in S_{d}(\Gamma_{0}(N))$ is a cusp form of weight $d$ and level $N$.
In the case of toric cycles $C_{\chi}$ attached to isotropic $\mathbb{Q}$-tori, as considered in [Bra23], the lift is not a cusp form in general. In fact, for suitable lattice in $V$ the constant term of $\Theta_{\varphi}(\tau,C_{\chi})$ is a partial Hecke $L$-function of $\chi$.
@@
for some $\alpha\in F^{\times}$. Moreover, we have $\operatorname{T}(\mathbb{Q})\simeq E^{1}$ where $E^{1}$ are the elements $x$ in $E$ of norm $x\epsilon(x)=1$. The quadratic extension $E/F$ is a product of quadratic extension $E_{i}/F_{i}$ where $F_{i}$ is a field. If $E_{i}=F_{i}\times F_{i}$ is split, the involution permutes the two factors and $E_{i}^{1}\simeq F_{i}^{\times}$. This is the case considered in [Bra23]. On the other hand, when $E_{i}/F_{i}$ is a field extension, the involution $\epsilon$ is the Galois involution $\textrm{Gal}(E_{i}/F_{i})$. This is the case we want to consider in this paper. The assumption that $\operatorname{T}$ is $\mathbb{Q}$-anisotropic implies that none of the factors $E_{i}$ is split. On the other hand, the assumption that the real rank of $\operatorname{T}$ is maximal implies that the fields $F_{i}$ are totally real.
@@
$\Theta_{\varphi}(\tau,C_{\chi})=\sum_{n=1}^{\infty}\langle C_{\chi},C_{n}(\varphi)\rangle q^{n}=\Theta_{\varphi}(\tau,\cdots,\tau,\chi)\in S_{d}(\Gamma_{0}(N))$ (1.13)
is the diagonal restriction of a Hilbert modular form $\Theta_{\varphi}(\tau_{1},\dots,\tau_{d},\chi)$ of parallel weight one for a congruence subgroup of $\operatorname{SL}_{2}(F)$.
@@ 1.3.  Spans of diagonal restrictions and toric cycles
Let us now add the conditions that $V$ is of dimension $p+q>4$, where $q$ is odd, and that the lattice $L$ is of level $1$ (unimodular). Let $S$ be the set of all characters $\chi$ that are as above, odd and of finite order. Let
@@
be the subspace of cusp forms spanned by the diagonal restrictions $\Theta_{\varphi}(C_{\chi})$. Several authors (including [Li17],[BL22],[KP19] and [Yan05] ) have considered spans of diagonal restrictions of Hilbert modular forms, mainly in the case of Hilbert-Eisenstein series.
@@
With the previous conditions on the signature of $V$, the adjoint of the Kudla-Millson lift is injective by a result of Bruinier-Funke [BF10]. In section 5 we deduce the following.
@@ 1.4.  Examples
Let us first consider the étale algebra $E=\mathbb{Q}(\sqrt{D})\times\mathbb{Q}(\sqrt{D})$ with involution being the Galois involution in both factors. After twisting by suitable characters, the toric cycle attached to $\operatorname{T}(\mathbb{Q})\simeq E^{1}$ is $\gamma_{\psi}\times\gamma_{\psi^{\prime}}$, where $\gamma_{\psi}$ and $\gamma_{\psi^{\prime}}$ are both attached to the same real quadratic field $K=\mathbb{Q}(\sqrt{D})$, where $p$ is split. This setting was considered by Darmon-Harris-Rotger-Venkatesh in [Dar+22], where they show that the generating series
@@
[…]
@@
Similarly, we can consider a biquadratic field $E=\mathbb{Q}(\sqrt{D_{1}},\sqrt{D_{2}})$ with the involution that sends $\sqrt{D_{i}}$ to $-\sqrt{D_{i}}$. The corresponding seesaw
[…]
@@ Acknowledgments.
We thank Henri Darmon for suggesting to look at these seesaws, Pierre Charollois and Luis Garcia for helpful comments.
@@
In this section, we review the relations between algebraic tori in orthogonal groups and étale algebras with involutions, as explained in [BCKM03], [BF14] and [DGL23].
@@
    the involution $\epsilon$ preserves the factor $R$ and ${\left.\kern-1.2ptR\vphantom{\big{|}}\right|_{\epsilon}}\neq\mathrm{id}$,
  • $ii)$
    the involution $\epsilon$ does not preserve the factor $R$. In that case there is another factor $R^{\prime}$ such that $R^{\prime}=\epsilon(R)\simeq R$,
@@
Hence we can write $E$ as a sum $E=E_{1}\times\cdots\times E_{r}$ of $\epsilon$-invariant subalgebras, where $E_{i}=R$ is a field in case $i)$ and $iii)$, and $E_{i}=R\times R^{\prime}$ is a product of fields in case $ii)$. The fixed algebra $F$ is then the sum $F=F_{1}\times\cdots\times F_{r}$ where $F_{i}=E_{i}^{\epsilon}$ is a number field. From now on suppose furthermore that $F$ has degree $\frac{1}{2}[E\colon\mathbb{Q}]$ over $\mathbb{Q}$. In that case we only have case $i)$ or $ii)$ and we deduce the following.
@@
There is an element $\delta\in F^{\times}$ such that $E$ is isomorphic to $F[\theta]/(\theta^{2}-\delta)$ and the involution $\epsilon$ sends $\theta$ to $-\theta$.
@@
For the case $i)$ the field extension $E_{i}/F_{i}$ is of degree $2$. Hence we have $E_{i}\simeq F_{i}[\theta_{i}]/(\theta_{i}^{2}-\delta_{i})$ for some $\delta_{i}$ in $F_{i}^{\times}\setminus(F_{i}^{\times})^{2}$. In case $ii)$ note we can identify $E_{i}\simeq F_{i}\times F_{i}$. Then we can take $\delta_{i}=1$ so that we get an isomorphism
@@
It is preserved by the elements of norm $1$
@@
Suppose that $F$ is a field and $E/F$ is an étale algebra. As mentionned in the introduction, the case where $E=F\times F$ is the split algebra will be excluded, hence we will restrict ourselves to the case where $E=F(\theta)$ is a quadratic field extension of $F$. In order to work with orthogonal groups instead of unitary groups, we view $E$ as an $F$-vector space and let $Q_{\alpha}$ be the quadratic form obtained by composing the hermitian form with the trace
@@
The parameter $\alpha$ is irrelevant here so let us assume $\alpha=1$. With respect to the basis $F$-basis $\{1,\theta\}$ of $E$, the quadratic form $Q_{\alpha}$ has Gram matrix $\operatorname{diag}(1,-\delta)$. It is clear that $E^{1}\subset\operatorname{SO}_{E}(F)$. On the other hand, for a matrix $g=\begin{pmatrix}a&b\\ c&d\end{pmatrix}$ the condition $g^{T}\begin{pmatrix}1&0\\ 0&-\delta\end{pmatrix}g=\begin{pmatrix}1&0\\ 0&-\delta\end{pmatrix}$ implies $a^{2}-c^{2}\delta=1$, $b^{2}-d^{2}\delta=-\delta$ and $ab=cd\delta.$ By multiplying the last equation by $d$ on both sides and using that $ad-bc=1$ we find that $b=c\delta$. Using the last equation again, we find that $ab=db$. If $b\neq 0$, this implies $a=d$. If $b=0$ one can easily check that $g=\pm\mathbf{1}_{2}$, hence $a=d=\pm 1$.
@@
$\operatorname{SO}_{E}(F)$⟶$\lhook\joinrel\longrightarrow\operatorname{SO}_{V}(\mathbb{Q}).⸦$ (2.14)
@@
Let $w$ be a non-archimedean place of $F_{i}$, where $E_{i}=F_{i}(\theta_{i})$. Then, either $E_{w}=F_{w}(\theta_{i})$ is non-split and the involution sends $\theta_{i}$ to $-\theta_{i}$, or $E_{w}=F_{w}\times F_{w}$ is split and the involution permutes the two factors. At a non-split place $w$ we have
@@
$S_{1}$$\coloneqq\{\ \textrm{real embeddings of $F$ that extend to real embeddings of $E$}\},$
$S_{2}$$\coloneqq\{\ \textrm{real embeddings of $F$ that extend to complex embeddings of $E$}\},$ (2.16)
@@
Let $(V,Q)$ be a non-degenerate quadratic $\mathbb{Q}$-space of dimension $2d$ with orthogonal group $\operatorname{SO}_{V}$. An algebraic $\mathbb{Q}$-group $\operatorname{T}\subset\operatorname{SO}_{V}$ is a $\mathbb{Q}$-torus of rank $a$ if $\operatorname{T}(\overline{\mathbb{Q}})\simeq\overline{\mathbb{Q}}^{a}$. If $K$ is a field extension of $\mathbb{Q}$, we say that $\operatorname{T}$ is $K$-split if $\operatorname{T}(K)\simeq K^{a}$. A torus $\operatorname{T}$ in the orthogonal group $\operatorname{SO}_{V}$ is maximal if its rank is equal to $\dim(V)$. The dimension of maximal $K$-split torus $\operatorname{T}^{\prime}\subset\operatorname{T}$ is called the $K$-rank of $\operatorname{T}$. The orthogonal group $\operatorname{SO}_{V}$ is of $\mathbb{R}$-rank $q$ (since we assume that $p\geq q$). Hence, if $\operatorname{T}$ is of maximal $\mathbb{R}$-rank, then the $\mathbb{R}$-rank of $\operatorname{T}$ is $q$.

### v4

# Romain Branchereau [^1] McGill, Burnside Hall 805 Sherbrooke Street West Montreal, Quebec H3A 0B9

@@ 1.1. Intersection numbers of geodesics on modular curves
[…]
@@
If two geodesics $\gamma_{1}$ and $\gamma_{2}$ in $Y_{0}(p)$ intersect transversely and in a compact set then $\langle\gamma_{1},\gamma_{2}\rangle=\sum_{z\in\gamma_{1}\cap\gamma_{2}}\pm 1$ is the topological intersection number, where $\pm 1$ depends on the local orientation at the intersection point. Darmon-Pozzi-Vonk prove the following in [10, Theorem. A].
@@
of weight $2$ and level $\Gamma_{0}(p)$ is the diagonal restriction of a $p$-stabilized Hilbert-Eisenstein series $E^{(p)}(\psi,\tau,\tau)$ for a subgroup $\operatorname{SL}_{2}(\mathbb{Q}(\sqrt{D}))$.
In [4], we showed how to recover (and generalize) the theorem of Darmon-Pozzi-Vonk by using the Kudla-Millson lift. The idea was to consider the embedding $F^{\times}\subset\operatorname{SO}(d,d)$ of a totally real field $F$ of degree $d$, which gives a non-compact cycle on the locally symmetric space of $\operatorname{SO}(d,d)$. The (regularized) integral of the Kudla-Millson theta lift over this cycle gives a generating series of intersection numbers generalizing the right handside of (1.4). On the other hand, the relation to a Hilbert-Eisenstein series follows from a Siegel-Weil formula and the seesaw
$\operatorname{SO}(d,d)$$\operatorname{SL}_{2}(F)$$F^{\times}$$\operatorname{SL}_{2}(\mathbb{Q}).$}
In the case where $F$ is a real quadratic field, this yields the theorem of Darmon-Pozzi-Vonk above.
@@
Another closely related seesaw appears in [2]
$\operatorname{GL}_{d}(\mathbb{Q})$$\operatorname{GL}_{2}(F)$$F^{\times}$$\operatorname{GL}_{2}(\mathbb{Q}),$}
where the right hand side is the diagonal restriction of (the same) Hilbert-Eisenstein series, and the left hand side is the evaluation of an Eisenstein class on the same cycle as above. The two seesaws are related by an embedding $\operatorname{GL}_{d}(\mathbb{Q})$ in $\operatorname{SO}(d,d)$, and the Bergeron-Charollois-Garcia lift is closely related to the Kudla-Millson lift via the Mathai-Quillen form. We hope to explain this in more detail in future work.
@@ 1.2. Main result
In this paper, we replace the torus $F^{\times}\subset\operatorname{SO}(d,d)$ by a more general anisotropic maximal $\mathbb{Q}$-torus, of maximal $\mathbb{R}$-rank in an orthogonal group $\operatorname{SO}(p,q)$ of signature $(p,q)$ with $p\geq q>0$. We restrict ourselves to anisotropic tori, to avoid having to deal with the regularization of the integral as in [4]. However, one can combine the result presented here with the result of loc. cit. to consider any maximal $\mathbb{Q}$-torus of maximal $\mathbb{R}$-rank, not necessarily anisotropic.
@@
The study of cycles attached to algebraic tori in orthogonal groups is also motivated by the recent work of Darmon-Gehrmann-Linowski. In [8], the authors define rigid meromorphic cocycles in arbitrary signature, generalizing the rigid meromorphic cocycles in signature $(2,2)$ studied in [10]. They conjecture that evaluated on toric cycles, these cocycles should be algebraic. The seesaws presented here could be relevant.
@@
Let $\varphi_{f}\in\mathcal{S}(V_{\mathbb{A}_{f}})$ be a finite Schwartz function preserved by an open compact subgroup $K_{f}\subset\operatorname{SO}_{V}(\mathbb{A}_{f})$. We set $K=K_{\infty}K_{f}$ and consider the adelic space
@@
In [14, 15], Kudla and Millson construct an element
$\Theta^{V}_{\varphi}\in H^{q}(Y)\otimes M_{d}(\Gamma^{\prime}_{\mathbb{Q}})$ (1.6)
that realizes a lift from the homology to the space of modular forms of weight $d=\frac{p+q}{2}$ for $\Gamma^{\prime}_{\mathbb{Q}}\subset\operatorname{SL}_{2}(\mathbb{Q})$ depending on the choice of the Schwartz function. As for the modular curve, there is a homological pairing
$\langle-,-\rangle\colon H_{q}(Y,\mathbb{Z})\times{H_{pq-q}(Y,\partial Y,\mathbb{Z})}\longrightarrow\mathbb{Z}.$ (1.7)
If the homology classes are represented by two smooth immersed submanifolds that intersect transversely and in a compact set, then the intersection number $\langle C_{1},C_{2}\rangle$ is the signed intersection number, as for geodesics. The main feature of the Kudla-Millson lift is that for a cycle $C$ in $H_{q}(Y)$ it has the Fourier expansion
$\Theta^{V}_{\varphi}(\tau,C)\coloneqq\int_{C}\Theta^{V}_{\varphi}(\tau)=c^{0}_{\varphi}(C)+\sum_{n\in\mathbb{Q}_{>0}}\langle C,C_{n}(\varphi)\rangle q^{n},$ (1.8)
@@
[…]
@@
[…]
@@
where $\Lambda\coloneqq\operatorname{T}(\mathbb{Q})^{+}\cap K_{T,f}$. The quotient is compact since $\operatorname{T}$ is anisotropic. The embedding of $\operatorname{T}(\mathbb{A})$ in $\operatorname{SO}_{V}(\mathbb{A})$ induces a map $Y_{T}\longrightarrow Y$, and the image of the connected component $\Lambda\backslash\mathbb{R}_{>0}^{q}$ associated to $I\in\mathcal{C}_{T}$ defines a cycle $C_{T,I}\in\mathcal{Z}_{q}(Y)$. Let us now suppose that $\chi_{f}$ is trivial on $K_{T,f}$. We can see the finite part $\chi_{f}$ as a function on $\chi_{f}\colon\mathcal{C}_{T}\longrightarrow\mathbb{C}^{\times}$ and define the cycle
@@
By the work of Kudla-Millson, it is known that if $q$ is odd, then $\Theta^{V}_{\varphi}(\tau,C)$ is a cusp form for any cycle $C$; see [17, Theorem. 2]. We show that when we restrict to the cycles $C_{\chi}$, then the lift is always a cusp form.
@@
Let $V$ be an even dimensional quadratic space of signature $(p,q)$ with $p\geq q>0$ and let $C_{\chi}$ be the cycle attached to an anisotropic maximal $\mathbb{Q}$-torus of maximal real rank $q$. Then $\Theta^{V}_{\varphi}(\tau,C_{\chi})\in S_{d}(\Gamma^{\prime}_{\mathbb{Q}})$ is a cusp form of weight $d$ for $\Gamma^{\prime}_{\mathbb{Q}}\subset\operatorname{SL}_{2}(\mathbb{Q})$.
In the case of toric cycles $C_{\chi}$ attached to isotropic $\mathbb{Q}$-tori, as considered in [4], the lift is not a cusp form in general. In fact, for suitable lattice in $V$ the constant term of $\Theta_{\varphi}(\tau,C_{\chi})$ is a partial Hecke $L$-function of $\chi$.
@@
for some $\alpha\in F^{\times}$. Moreover, we have $\operatorname{T}(\mathbb{Q})\simeq E^{1}$ where $E^{1}$ are the elements $x$ in $E$ of norm $x\epsilon(x)=1$. The quadratic extension $E/F$ is a product of quadratic extension $E_{i}/F_{i}$ where $F_{i}$ is a field. If $E_{i}=F_{i}\times F_{i}$ is split, the involution permutes the two factors and $E_{i}^{1}\simeq F_{i}^{\times}$. This is the case considered in [4]. On the other hand, when $E_{i}/F_{i}$ is a field extension, the involution $\epsilon$ is the Galois involution $\textrm{Gal}(E_{i}/F_{i})$. This is the case we want to consider in this paper. The assumption that $\operatorname{T}$ is $\mathbb{Q}$-anisotropic implies that none of the factors $E_{i}$ is split. On the other hand, the assumption that the real rank of $\operatorname{T}$ is maximal implies that the fields $F_{i}$ are totally real.
@@
$\Theta_{\varphi}(\tau,C_{\chi})=\sum_{n\in\mathbb{Q}_{>0}}\langle C_{\chi},C_{n}(\varphi)\rangle q^{n}=\Theta_{\varphi}(\tau,\cdots,\tau,\chi)\in S_{d}(\Gamma^{\prime}_{\mathbb{Q}})$ (1.13)
is the diagonal restriction of a Hilbert modular form $\Theta_{\varphi}(\tau_{1},\dots,\tau_{d},\chi)$ of parallel weight one for some subgroup of $\operatorname{SL}_{2}(F)$.
@@ 1.3. Spans of diagonal restrictions and toric cycles
Let us now add the conditions that $V$ is of dimension $p+q>4$, where $q$ is odd. Let $L\subset V$ be an even unimodular lattice and $\varphi_{f}=\mathbf{1}_{\widehat{L}}$ where $\widehat{L}=L\otimes\widehat{\mathbb{Z}}$. Let $S$ be the set of all characters $\chi$ that are as above, odd and of finite order. Let
@@
be the subspace of cusp forms spanned by the diagonal restrictions $\Theta_{\varphi}(C_{\chi})$. Several authors (including [19],[3],[12] and [23] ) have considered spans of diagonal restrictions of Hilbert modular forms, mainly in the case of Hilbert-Eisenstein series.
@@
With the previous conditions on the signature of $V$, the adjoint of the Kudla-Millson lift is injective by a result of Bruinier-Funke [6]. In section 5 we deduce the following.
@@
Remark 1.4.
Corollary 1.2.1 also holds for non-unimodular lattices, provided that the Kudla-Millson lift is injective. In particular, one could use stronger injectivity results of the Kudla-Millson lift due to Zuffetti, Metzler and Kiefer [24, 20, 13], generalizing the result of Bruinier-Funke in the unimodular case.
@@ 1.4. Examples
Let us first consider the étale algebra $E=\mathbb{Q}(\sqrt{D})\times\mathbb{Q}(\sqrt{D})$ with involution being the Galois involution in both factors. After twisting by suitable characters, the toric cycle attached to $\operatorname{T}(\mathbb{Q})\simeq E^{1}$ is $\gamma_{\psi}\times\gamma_{\psi^{\prime}}$, where $\gamma_{\psi}$ and $\gamma_{\psi^{\prime}}$ are both attached to the same real quadratic field $K=\mathbb{Q}(\sqrt{D})$, where $p$ is split. This setting was considered by Darmon-Harris-Rotger-Venkatesh in [9], where they show that the generating series
@@
$\operatorname{SO}(2,2)$$\operatorname{SL}_{2}(\mathbb{Q})\times\operatorname{SL}_{2}(\mathbb{Q})$$\mathbb{Q}(\sqrt{D})^{1}\times\mathbb{Q}(\sqrt{D})^{1}$$\operatorname{SL}_{2}(\mathbb{Q}),$}
@@
Similarly, we can consider a biquadratic field $E=\mathbb{Q}(\sqrt{D_{1}},\sqrt{D_{2}})$ with the involution that sends $\sqrt{D_{i}}$ to $\sqrt{D_{i}}$. The corresponding seesaw
$\operatorname{SO}(2,2)$$\operatorname{SL}_{2}(\mathbb{Q}(\sqrt{D_{1}D_{2}}))$$\mathbb{Q}(\sqrt{D_{1}},\sqrt{D_{2}})^{1}$$\operatorname{SL}_{2}(\mathbb{Q}),$}
@@ Acknowledgments.
We thank Henri Darmon for suggesting to look at these seesaws, Pierre Charollois, Luis Garcia, Håvard Damm-Johnsen and the referees for helpful comments.
@@
In this section, we review the relations between algebraic tori in orthogonal groups and étale algebras with involutions, as explained in [7], [1] and [8].
@@
    the involution $\epsilon$ preserves each factor $R$ and ${\left.\kern-1.2pt\epsilon\vphantom{\big|}\right|_{R}}\neq\mathrm{id}$,
  • ii)ii)
    the involution $\epsilon$ does not preserve one of the factors $R$. In that case there is another factor $R^{\prime}$ such that $R^{\prime}=\epsilon(R)\simeq R$,
@@
Hence we can write $E$ as a sum $E=E_{1}\times\cdots\times E_{r}$ of $\epsilon$-invariant subalgebras, where $E_{i}=R$ is a field in case $i)$ and $iii)$, and $E_{i}=R\times R^{\prime}$ is a product of fields in case ii)ii). The fixed algebra $F$ is then the sum $F=F_{1}\times\cdots\times F_{r}$ where $F_{i}=E_{i}^{\epsilon}$ is a number field. From now on suppose furthermore that $F$ has degree $\frac{1}{2}[E\colon\mathbb{Q}$] over $\mathbb{Q}$. In that case we only have case $i)$ or ii)ii) and we deduce the following.
@@
There is an element $\delta\in F^{\times}$ such that $E$ is isomorphic to $F[\theta]/(\theta^{2}-\delta)$ and the involution $\epsilon$ sends $\theta$ to $\theta$.
@@
For the case $i)$ the field extension $E_{i}/F_{i}$ is of degree $2$. Hence we have $E_{i}\simeq F_{i}[\theta_{i}]/(\theta_{i}^{2}-\delta_{i})$ for some $\delta_{i}$ in $F_{i}^{\times}\setminus(F_{i}^{\times})^{2}$. In case ii)ii) note we can identify $E_{i}\simeq F_{i}\times F_{i}$. Then we can take $\delta_{i}=1$ so that we get an isomorphism
@@
It is preserved by the elements of norm 11
@@
Suppose that $F$ is a field and $E/F$ is an étale algebra. As mentioned in the introduction, the case where $E=F\times F$ is the split algebra will be excluded, hence we will restrict ourselves to the case where $E=F(\theta)$ is a quadratic field extension of $F$. In order to work with orthogonal groups instead of unitary groups, we view $E$ as an $F$-vector space and let $Q_{\alpha}$ be the quadratic form obtained by composing the hermitian form with the trace
@@
The parameter $\alpha$ is irrelevant here so let us assume $\alpha=1$. With respect to the basis $F$-basis $\{1,\theta\}$ of $E$, the quadratic form $Q_{\alpha}$ has Gram matrix $\operatorname{diag}(1,-\delta)$. It is clear that $E^{1}\subset\operatorname{SO}_{E}(F)$. On the other hand, for a matrix $g=\begin{pmatrix}a&b\\ c&d\end{pmatrix}$ the condition $g^{T}\begin{pmatrix}1&0\\ 0&-\delta\end{pmatrix}g=\begin{pmatrix}1&0\\ 0&-\delta\end{pmatrix}$ implies $a^{2}-c^{2}\delta=1$, $b^{2}-d^{2}\delta=-\delta$ and $ab=cd\delta.$ By multiplying the last equation by $d$ on both sides and using that adbc=1ad-bc=1 we find that $b=c\delta$. Using the last equation again, we find that $ab=db$. If $b\neq 0$, this implies $a=d$. If $b=0$ one can easily check that $g=\pm\mathbf{1}_{2}$, hence $a=d=\pm 1$.
@@
$\operatorname{SO}_{E}(F)$ $\lhook\joinrel\longrightarrow\operatorname{SO}_{V}(\mathbb{Q}).$ (2.10)
@@
Let $w$ be a non-archimedean place of $F_{i}$, where $E_{i}=F_{i}(\theta_{i})$. Then, either $E_{w}=F_{w}(\theta_{i})$ is non-split and the involution sends $\theta_{i}$ to $\theta_{i}$, or $E_{w}=F_{w}\times F_{w}$ is split and the involution permutes the two factors. At a non-split place $w$ we have
@@
Let $(V,Q)$ be a non-degenerate quadratic $\mathbb{Q}$-space of dimension $2d$ with orthogonal group $\operatorname{SO}_{V}$. An algebraic $\mathbb{Q}$-group $\operatorname{T}\subset\operatorname{SO}_{V}$ is a $\mathbb{Q}$-torus of rank $a$ if $\operatorname{T}(\overline{\mathbb{Q}})\simeq\overline{\mathbb{Q}}^{a}$. If $K$ is a field extension of $\mathbb{Q}$, we say that $\operatorname{T}$ is $K$-split if $\operatorname{T}(K)\simeq K^{a}$. A torus $\operatorname{T}$ is maximal if it is not contained in a torus of larger rank. The rank of a maximal $K$-split torus $\operatorname{T}^{\prime}\subset\operatorname{T}$ is called the $K$-rank of $\operatorname{T}$. The orthogonal group $\operatorname{SO}_{V}$ is of $\mathbb{R}$-rank $q$ (since we assume that $p\geq q$). Hence, if $\operatorname{T}$ is of maximal $\mathbb{R}$-rank, then the $\mathbb{R}$-rank of $\operatorname{T}$ is $q$.